OpenStudy (mikee):

A brownie recipe asks for one and one third times as much sugar as chocolate chips. If three and one half cups of sugar is used, what quantity of chocolate chips would then be needed, according to the recipe?

OpenStudy (dumbcow):

1 and 1/3 = 4/3 3 and 1/2 = 7/2 ratio: \[\frac{\frac{4}{3}}{1} = \frac{\frac{7}{2}}{x}\] \[x = \frac{7}{2}*\frac{3}{4} = \frac{21}{8}\] 21/8 = 2 and 5/8

OpenStudy (mikee):

Ty